# A Mind Cannot Be Smeared Across Time

### v1

@@ 4.3 The algebraic core
The following commutation and non-commutation facts are elementary, but spelling them out in Stack-Theory form turns the Temporal Gap into a concrete architectural question.
@@
$\Box_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Box_{\Delta}l).$
Proof.
Fix any window $\sigma\in\Phi^{\Delta}$. Then
$\sigma\in\mathrm{T}(\Box_{\Delta}l)$ $\iff\forall p\in l,\ \sigma\in\Box_{\Delta}p$
$\iff\forall p\in l,\ \forall k\leq\Delta,\ \sigma(k)\in p$
$\iff\forall k\leq\Delta,\ \sigma(k)\in\bigcap_{p\in l}p$
$\iff\sigma\in\Box_{\Delta}\mathrm{T}(l).$
Proof intuition.
$\Box_{\Delta}$ is “for all positions in the window”. Since $\forall$ distributes over conjunction, $\Box_{\Delta}$ commutes with $\wedge$.
@@
If $\Delta=0$ (a window is a single instant) or $\lvert l\rvert\leq 1$ (no nontrivial conjunction), then
$\Diamond_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Diamond_{\Delta}l).$
Thus strict separation requires $\Delta\geq 1$ and at least two ingredients.
@@
$\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}l),$
$\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
Thus Theorem 3 is exactly the claim that ingredient-wise satisfaction is strictly weaker than co-instantiation in general.
4.4 When does $\Diamond$ commute with conjunction?
Definition 15 (Within-window persistence along a trajectory).
Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau:\mathbb{N}\to\Phi$. A program $p\subseteq\Phi$ is $(\tau,\Delta)$-persistent if for all $t\in\mathbb{N}$ and all indices $k\leq k^{\prime}\leq\Delta$,
$\tau(t+k)\in p\ \Longrightarrow\ \tau(t+k^{\prime})\in p.$
Intuition.
Along the trajectory, if $p$ becomes true at some point inside a $\Delta$ window, then $p$ stays true for the rest of that same window.
Proposition 2 (A sufficient condition for co-instantiation along a trajectory).
Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau$. Let $l\in\mathcal{L}_{\mathfrak{v}}$ be a statement such that every ingredient program $p\in l$ is $(\tau,\Delta)$-persistent. Then for all $t$,
$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)\quad\Longrightarrow\quad\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
Proof.
Assume $\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)$. Then for each $p\in l$ there exists $k_{p}\leq\Delta$ with $\tau(t+k_{p})\in p$. Let $k^{\ast}:=\max_{p\in l}k_{p}$. By $(\tau,\Delta)$-persistence of each $p$, from $\tau(t+k_{p})\in p$ and $k_{p}\leq k^{\ast}\leq\Delta$ I obtain $\tau(t+k^{\ast})\in p$ for all $p\in l$. Hence $\tau(t+k^{\ast})\in\bigcap_{p\in l}p=\mathrm{T}(l)$, so $\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l)$. ∎
Proof intuition.
With persistence, once each ingredient appears it remains true. Choosing the latest appearance time yields a single instant where all ingredients overlap.
4.5 Finite-horizon temporal semantics
Finite-horizon (history-dependent) temporal semantics can itself be treated as an abstraction layer. A horizon-$\Delta$ pathwise abstractor can be compiled into a pointwise abstractor on an augmented state, so Stack-Theory tools apply without modifying the underlying semantics.
Definition 16 (Pointwise and horizon-$\Delta$ pathwise abstractors).
Let $\Psi$ be any higher-level state space. A pointwise abstractor is a function $A:\Phi\to\Psi$. A pathwise abstractor of horizon $\Delta$ is a function $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$.
Intuition.
A pointwise abstractor looks only at the current state. A horizon $\Delta$ pathwise abstractor looks at the recent $\Delta{+}1$ states, so it can use short term history.
Theorem 4 (History-state (state augmentation) compilation).
Fix $\Delta\in\mathbb{N}$ and let $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$ be a pathwise abstractor. Define the augmented environment $\widetilde{\Phi}:=\Phi^{\Delta}$ and the pointwise abstractor $\widetilde{A}:\widetilde{\Phi}\to\Psi$ by $\widetilde{A}(\sigma):=A^{(\Delta)}(\sigma)$. Then for every objective trajectory $\tau:\mathbb{N}\to\Phi$ and every $t$, with augmented trajectory $\widetilde{\tau}(t):=\tau^{\Delta,1}(t)\in\widetilde{\Phi}$ I have
$\widetilde{A}(\widetilde{\tau}(t))=A^{(\Delta)}(\tau^{\Delta,1}(t)).$
Proof.
Immediate from the definitions of $\widetilde{A}$ and $\widetilde{\tau}$. ∎
Proof intuition.
Encode the last $\Delta$ microstates into an augmented state so any horizon-$\Delta$ (history-dependent) abstractor becomes pointwise on the augmented state.
5 StrongSync and WeakSync Postulates
The above algebra does not decide which window semantics corresponds to phenomenality. It clarifies what must be decided. I’ll now state StrongSync and WeakSync in terms of a higher-layer moment statement and its base-layer grounding across a stack state.
@@ 5 StrongSync and WeakSync Postulates
Let $\mathrm{PhenReal}(l^{m},\tau,t)$ mean “moment statement $l^{m}$ is phenomenally realised at (higher-layer) time $t$ along objective trajectory $\tau$.” I treat $\mathrm{PhenReal}$ as primitive, constrained by StrongSync/WeakSync.
@@
$\mathrm{PhenReal}$ names the claim that moment statement $l^{m}$ is experienced at subjective time $t$ along the run $\tau$. Here $l^{m}$ can include first, second, and third order selves and causal identities for objects and properties [undeff].
@@
Intuition.
$g^{0}$ is the low level conjunction that corresponds to the higher level moment statement $l^{m}$. It is built by expanding $l^{m}$ down to the base layer. It is true in exactly the same situations as $l^{m}$.
Definition 19 (StrongSync (objective co-instantiation required)).
Fix a windowing map for the phenomenal layer, typically the sliding window $\mathrm{W}_{\Delta,s}$ of horizon $\Delta$ and stride $s$ (Section 4.2). StrongSync postulates that for all $l^{m},\tau,t$,
$\mathrm{PhenReal}(l^{m},\tau,t)$ $\Rightarrow\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
$\text{(equivalently }\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\text{).}$
@@
StrongSync says that if the moment is experienced, then within the window there is at least one objective time step where all grounded ingredients are true together.
Definition 20 (WeakSync (objective smearing permitted)).
Fix the same windowing map $\mathrm{W}_{\Delta,s}$ as in Definition 19. WeakSync postulates the following.
@@
    $\mathrm{PhenReal}(l^{m},\tau,t)\Rightarrow\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
    $\text{(equivalently}\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\text{).}$
@@
    $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)\wedge\neg\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t),$
    equivalently $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\wedge\tau^{\Delta,s}(t)\notin\Diamond_{\Delta}\mathrm{T}(g^{0})$.
In plain English this allows a moment to be experienced without any instant realising the full grounded conjunction.
@@
WeakSync says that if the moment is experienced, then each grounded ingredient must occur somewhere in the window. It allows that there may be no single objective time step where they are all true together.
@@ 6 Architectural Consequences
A polycomputer simultaneously executes multiple functions using the same parts or physical substrate [undefk]. A liquid brain such as an ant colony does not rely on a persistent structure to compute, but instead computes through movement [undefx]. In contrast, a “solid brain” is one that has a persistent structure that can support, for example, a bio-electric abstraction layer with synchronous broadcast [undeff] (like a network of telegraph cables). I now translate “single-thread sequential emulation” versus “synchronous, solid-brained, polycomputational multi-contributor realisation” into a minimal architectural invariant with a bound on simultaneous causal contribution. The Temporal Gap determines whether that bound is consciousness-relevant.
@@ 6.1 Concurrency capacity as an invariant
Fix $n\geq 2$ potential contributors and write $n]:=\{1,\dots,n\}$. Define a base environment of microstates
$\Phi:=X\times\mathcal{P}([n]),$
where $x\in X$ is a “data” component and $A\subseteq[n]$ is the set of currently active contributors at that objective instant. For each $i\in[n]$ define the program
$p_{i}:=\{(x,A)\in\Phi:i\in A\}.$
Let $l_{n}:=\{p_{1},\dots,p_{n}\}$ and note $\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$.
@@
The contributor model records which contributors are active at each time step. A conjunction that mentions many contributors can only be true when all of those contributors are active at the same time step.
@@
An architecture has concurrency capacity $c\in\{1,\dots,n\}$ if its admissible trajectories $\tau:\mathbb{N}\to\Phi$ satisfy (writing $\tau(t)=(x_{t},A_{t})$ for the data and contributor-set components)
$\lvert A_{t}\rvert\leq c\quad\text{for all objective times }t.$
The case $c=1$ formalises a sequential single-thread update model (one active contributor per tick). The case $c\geq n$ formalises synchronous polycomputing at the relevant scale (all contributors may be active together).
@@
Capacity $c$ is the maximum number of contributors that can be active at one time step. If the grounded content needs more than $c$ contributors active together, then StrongSync cannot hold on that architecture.
@@
Fix $n\geq 2$ and content statement $l_{n}$. If an architecture has concurrency capacity $c<n$, then along every admissible trajectory $\tau$ and for every horizon $\Delta$ I have
$\tau^{\Delta,1}(t)\notin\Diamond_{\Delta}\mathrm{T}(l_{n})\quad\text{for all }t.$
However, if the architecture is such that every contributor is activated at least once within some horizon-$\Delta$ window (i.e. for some $t$ and all $i$ there exists $k_{i}\leq\Delta$ with $i\in A_{t+k_{i}}$), then
$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l_{n}).$
Thus the architecture admits ingredient-wise satisfaction but forbids objective co-instantiation.
@@
If $c<n$, then at any time $t$ I have $\lvert A_{t}\rvert\leq c<n$, so $A_{t}\neq[n]$. Hence $\tau(t)\notin\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$ for all $t$. Therefore no window contains an element of $\mathrm{T}(l_{n})$, so $\Diamond_{\Delta}\mathrm{T}(l_{n})$ is never satisfied. For the second claim, if each $i$ appears in $A_{t+k_{i}}$ for some $k_{i}\leq\Delta$, then the window $\sigma=\tau^{\Delta,1}(t)$ satisfies $\sigma\in\Diamond_{\Delta}p_{i}$ for all $i$, hence $\sigma\in\bigcap_{i}\Diamond_{\Delta}p_{i}=\mathrm{T}(\Diamond_{\Delta}l_{n})$. ∎
@@
Corollary 1 (Single-thread CPU vs synchronous polycomputer (formal)).
Let a moment statement $l^{m}$ have base grounding $g^{0}$ whose ingredients include a conjunctive substatement equivalent to $l_{n}$ for some $n\geq 2$. Fix the phenomenal windowing map $\mathrm{W}_{\Delta,s}$ used in StrongSync/WeakSync.
@@
    Under StrongSync, any sequential-update architecture with capacity $c<n$ cannot phenomenally realise $l^{m}$, because $\Diamond_{\Delta}\mathrm{T}(g^{0})$ fails for the relevant grounded substatement (Theorem 5).
@@
    Under WeakSync, the same sequential architecture remains a candidate in principle, because $\mathrm{T}(\Diamond_{\Delta}g^{0})$ may hold without $\Diamond_{\Delta}\mathrm{T}(g^{0})$.
@@
    Any synchronous architecture with capacity $c\geq n$ can satisfy $\Diamond_{\Delta}\mathrm{T}(g^{0})$ whenever it activates the required contributors simultaneously, hence can host the moment under both StrongSync and WeakSync.
@@
My valence-first application of Stack Theory treats certain phenomenal configurations as synchronised, causally efficacious conjunctions over many ingredients [undeff]. The capacity model above isolates one minimal implementation parameter, simultaneous contribution, that controls whether such conjunctions can be instantiated at any objective instant versus only “smeared” across a window.
7 Evidence for StrongSync
StrongSync and WeakSync predict different empirical signatures within a short subjective window. Either consciousness requires a window-local episode of coordinated co-engagement (objective co-instantiation), or it can tolerate mere ingredient-wise activation distributed across the window. Several classic paradigms are more naturally described by the coordination view. In masking, conscious perception has been reported to correlate with transient long-range gamma phase synchrony across cortical areas, even when local activity changes are similar across seen versus unseen stimuli [undeft]. In non-REM sleep, TMS-EEG responses can become strong yet local, failing to propagate, consistent with a breakdown of effective connectivity when consciousness fades [undefs]. More generally, perturbational measures such as PCI track level of consciousness across wake/sleep/anesthesia and disorders of consciousness, consistent with a requirement for integrated, differentiated system-level dynamics [undefm]. Taken together, these findings do not prove StrongSync, but they do make the existential permissiveness of WeakSync much less compelling. Conscious level appears to covary with the presence (and disappearance with the collapse) of a temporally coordinated, system-level regime. Hence when I draw architectural conclusions below, I treat StrongSync as the default working hypothesis. This step remains defeasible.

### v2

@@ 4.3 The Algebraic Core
For any statement $l\in\mathcal{L}_{\mathfrak{v}}$ and any $\Delta\in\mathbb{N}$, $\Box_{\Delta}\mathrm{T}(l)=\mathrm{T}(\Box_{\Delta}l)$.
Proof. See Supplementary Material. $\Box_{\Delta}$ is “for all positions”; since $\forall$ distributes over conjunction, $\Box_{\Delta}$ commutes with $\wedge$. ∎
@@
If $\Delta=0$ or $\lvert l\rvert\leq 1$, then $\Diamond_{\Delta}\mathrm{T}(l)=\mathrm{T}(\Diamond_{\Delta}l)$. Strict separation requires $\Delta\geq 1$ and at least two ingredients. A sufficient condition for commutation along a trajectory (within-window persistence) is given in the Supplementary Material.
@@
$\mathrm{Occur}_{\mathrm{W}_{\Delta}}(l,\tau,t)$ $\iff\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}l),$
$\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(l,\tau,t)$ $\iff\tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
Theorem 3 is exactly the claim that ingredient-wise satisfaction is strictly weaker than co-instantiation in general.
5 Chord and Arpeggio Postulates
The above algebra does not decide which window semantics corresponds to phenomenality. I now state Chord and Arpeggio in terms of occurrence and co-instantiation, with a higher-layer moment statement and its base-layer grounding.
@@ 5 Chord and Arpeggio Postulates
Let $\mathrm{PhenReal}(l^{m},\tau,t)$ mean “moment statement $l^{m}$ is phenomenally realised at (higher-layer) time $t$ along objective trajectory $\tau$.” I treat $\mathrm{PhenReal}$ as primitive, constrained by Chord/Arpeggio.
@@
$\mathrm{PhenReal}$ names the claim that moment $l^{m}$ is experienced at subjective time $t$ along run $\tau$. Here $l^{m}$ can include first, second, and third order selves and causal identities for objects and properties [7].
@@
Definition 16 (Chord (objective co-instantiation required)).
Fix a windowing map for the phenomenal layer, typically the sliding window $\mathrm{W}_{\Delta}$ of horizon $\Delta$ (Section 4.2). Chord postulates that for all $l^{m},\tau,t$,
$\mathrm{PhenReal}(l^{m},\tau,t)$ $\Rightarrow\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)$
$\text{(equivalently }\tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\text{).}$
@@
Chord says: if the moment is experienced, then within the window there is at least one objective time step where all grounded ingredients are true together.
Definition 17 (Arpeggio (objective smearing permitted)).
Fix the same windowing map $\mathrm{W}_{\Delta}$ as in Definition 16. Arpeggio postulates the following.
@@
    $\mathrm{PhenReal}(l^{m},\tau,t)\Rightarrow\mathrm{Occur}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)$
    $\text{(equivalently}\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\text{).}$
@@
    $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\mathrm{Occur}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)\wedge\neg\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t).$
This allows a moment to be experienced without any instant realising the full grounded conjunction.
@@
Arpeggio says: if the moment is experienced, each grounded ingredient must occur somewhere in the window, but there need not be a single time step where they are all true together.
@@ 6 Architectural Consequences
A polycomputer simultaneously executes multiple functions using the same parts or physical substrate [13]. A liquid brain such as an ant colony computes through movement rather than persistent structure [25]. In contrast, a “solid brain” has persistent structure supporting, e.g., a bio-electric abstraction layer with synchronous broadcast [7]. I now translate “single-thread sequential emulation” versus “synchronous polycomputational realisation” into a minimal architectural invariant bounding simultaneous causal contribution. The Temporal Gap asks whether that bound is consciousness-relevant (and whether conscious unity requires causal exchange between co-instantiated ingredients, explored in [10]).
@@ 6.1 Concurrency Capacity as an Invariant
Fix $n\geq 2$ potential contributors and write $n]:=\{1,\dots,n\}$. Define a base environment $\Phi:=X\times\mathcal{P}([n])$, where $x\in X$ is a “data” component and $A\subseteq[n]$ is the set of currently active contributors. For each $i\in[n]$ define $p_{i}:=\{(x,A)\in\Phi:i\in A\}$. Let $l_{n}:=\{p_{1},\dots,p_{n}\}$ and note $\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$.
@@
The contributor model records which contributors are active at each time step. A conjunction mentioning many contributors can only be true when all are active simultaneously.
@@
An architecture has concurrency capacity $c\in\{1,\dots,n\}$ if its admissible trajectories satisfy $\lvert A_{t}\rvert\leq c$ for all objective times $t$ (writing $\tau(t)=(x_{t},A_{t})$). The case $c=1$ formalises a sequential single-thread model; $c\geq n$ formalises synchronous polycomputing at the relevant scale.
@@
Capacity $c$ is the maximum number of contributors active at one time step. If the grounded content needs more than $c$ active together, co-instantiation cannot hold.
@@
Fix $n\geq 2$ and content statement $l_{n}$. If an architecture has concurrency capacity $c<n$, then along every admissible trajectory $\tau$ and for every horizon $\Delta$, $\tau_{\Delta}(t)\notin\Diamond_{\Delta}\mathrm{T}(l_{n})$ for all $t$. However, if every contributor is activated at least once within some horizon-$\Delta$ window, then $\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}l_{n})$. Thus the architecture admits ingredient-wise satisfaction but forbids objective co-instantiation.
@@
If $c<n$, then $\lvert A_{t}\rvert\leq c<n$ at every $t$, so $A_{t}\neq[n]$ and $\tau(t)\notin\mathrm{T}(l_{n})$ for all $t$. No window contains an element of $\mathrm{T}(l_{n})$, so $\Diamond_{\Delta}\mathrm{T}(l_{n})$ is never satisfied. For the second claim, if each $i$ appears in $A_{t+k_{i}}$ for some $k_{i}\leq\Delta$, then $\sigma=\tau_{\Delta}(t)$ satisfies $\sigma\in\Diamond_{\Delta}p_{i}$ for all $i$, hence $\sigma\in\mathrm{T}(\Diamond_{\Delta}l_{n})$. ∎
@@
Corollary 1 (Single-thread CPU vs synchronous polycomputer).
Let a moment statement $l^{m}$ have base grounding $g^{0}$ whose ingredients include a conjunctive substatement equivalent to $l_{n}$ for some $n\geq 2$. Fix the phenomenal windowing map $\mathrm{W}_{\Delta}$ used in Chord/Arpeggio.
@@
    Under Chord, any sequential-update architecture with capacity $c<n$ cannot phenomenally realise $l^{m}$ (Theorem 4).
@@
    Under Arpeggio, the same sequential architecture remains a candidate, because $\mathrm{T}(\Diamond_{\Delta}g^{0})$ may hold without $\Diamond_{\Delta}\mathrm{T}(g^{0})$.
@@
    Any synchronous architecture with capacity $c\geq n$ can satisfy $\Diamond_{\Delta}\mathrm{T}(g^{0})$ by activating the required contributors simultaneously, hence can host the moment under both postulates.
@@
My valence-first application of Stack Theory treats certain phenomenal configurations as synchronised, causally efficacious conjunctions over many ingredients [7]. The capacity model isolates one minimal implementation parameter—simultaneous contribution—controlling whether such conjunctions can be instantiated at any objective instant versus only “smeared” across a window.
7 Evidence for Chord
Chord and Arpeggio predict different empirical signatures within a short subjective window. Either consciousness requires window-local coordinated co-engagement, or it can tolerate mere ingredient-wise activation distributed across the window. Several classic paradigms are more naturally described by the coordination view. In masking, conscious perception correlates with transient long-range gamma phase synchrony across cortical areas, even when local activity is similar across seen versus unseen stimuli [21]. In non-REM sleep, TMS-EEG responses become strong yet local, failing to propagate, consistent with breakdown of effective connectivity when consciousness fades [20]. Perturbational measures such as PCI track level of consciousness across wake/sleep/anesthesia and disorders of consciousness, consistent with a requirement for integrated, differentiated dynamics [15]. These findings do not prove Chord, but they make the existential permissiveness of Arpeggio less compelling. Conscious level covaries with the presence of a temporally coordinated, system-level regime. I therefore treat Chord as the default working hypothesis; this step remains defeasible.